[figure: 9.7 cm]

EFFECT OF PRESSURE ON THE MAGNETIC PROPERTIES OF

RMn_{6}Ge_{6−x}Ga_{x} (R= Lu, Sc, Y) SOLID SOLUTIONS

R. Duraj^{a}, A. Szytuła^{b} and G. Venturini^{c} aInstitute of Physics, Technical University of Cracow Podchora̧żych 1, 30-084 Cracow, Poland

bInstitute of Physics, Jagiellonian University, Reymonta 4, 30-059 Cracow, Poland cLaboratoire de Chimie du Solide Mineral, Université Henri Poincaré -Nancy I 54506 Vandoeuvre les Nancy Cedex, France

We present results of the magnetic measurements of hexagonal HfFe_{6}Ge_{6}-type solid solutions RMn_{6}Ge_{6−x}Ga_{x} (R= Lu, Sc, Y; 0.1≤x≤1.0) in high pressure conditions. The changes in magnetic transition temperatures under external pressure up to 1.5GPa were determined by using the ac susceptibility method. The measurements of magnetization were carried out under isobaric conditions in a pulsed magnetic field at various temperatures. The pressure has almost no effect on the magnetic properties of scandium and itrium compounds and noticeable effect on the magnetic properties of the lutetium compound (x=0.7). Temperature of metamagnetic transition T_{1} strongly decrease with increasing pressure (dT_{1}/dP = -46K/GPa). The considerable temperature hysteresis of this transition was observed. Moreover, it seems that pressure and temperature act in opposite ways on the magnetic behaviour (considering that the F-AF transition is driven by a criterion based on the interatomic distances). The B-T curves show that, for a given temperature, the thereshold field increases with pressure (dB_{t}/dT = 1.5mT/K).

[figure: 13.4 cm]

Subject category :

5. Phase Transitions and Critical Phenomena

Presentation mode : poster

Corresponding author : R. Duraj

Address for correspondence : Dr R.Duraj Institute of Physics,Technical University of Cracow Podchora̧żych 1, 30-084 Cracow, Poland

Email address : puduraj@cyf-kr.edu.pl